OpenStudy (anonymous):

Let f(x)=log square root of x to base of 3. Show that f^-1(x)=3^2x. Write down the range of the inverse of f(x)

OpenStudy (anonymous):

\[f(x) = y = \log_3{\sqrt{x}}\] to find the inverse of a function, we would swap x and y in the function and the put it back in the form of f(x). giving us: \[x = \log_3{\sqrt{y}}\] for the general case: \[z = \log_b (d) \rightarrow d = b^{z}\] so in the inverse of the function \[\sqrt{y} = 3^x \rightarrow y = (3^x)^2 = 3^{2x}\] let me know if you have any questions ^_^

OpenStudy (anonymous):

so how would i find the range of the inverse?